Q 16 Question 16 Julia Penfield was working on a project and needed to update her...

Q 16

Question 16

Julia Penfield was working on a project and needed to update her boss on it’s CPI. If EV = 50,000, PV = 45,000, and AC = 41,000 what is the CPI?

Select one:

a. 1.21

b. 0.82

c. 1.16

d. 1.20

Q 17

Question 17

Kerri Lerner wanted to determine her project’s SPI. If EV = 42,000, PV = 40,000, and AC = 39,000, what is the SPI?

Select one:

a. 0.95

b. 0.98

c. 1.16

d. 1.05

Q 18

Question 18

If EV = 25,000, PV = 20,000, and AC = 29,000, what is the SV?

Select one:

a. -5,000

b. -1,000

c. 1,000

d. 5,000

Q 19

Question 19

If EV = 29,000, PV = 30,000, and AC = 25,000, what is the CV?

Select one:

a. -1,000

b. -4,000

c. 1,000

d. 4,000
Answer #1

Q 16) EV = 50000

AC = 41000

CPI = EV/AC = 50000/41000 = 1.21

Q17) EV = 42000

PV = 40000

SPI = EV/PV = 42000/40000 = 1.05

Q18) EV = 25000

PV = 20000

SV = EV-PV = 25000-20000 = 5000

Q19) EV = 29000

AC = 25000

CV = EV-AC = 29000-25000 = 4000